Assuming the earth to be a homogeneous sphere of radius R, its density in terms of G (constant of gravitation) and g (acceleration due to gravity on the surface of the earth)

Grade:11

1 Answers

Vikas TU
3 years ago
Using the relation,g = GM/R^2
Or M=gR^2/G−−−(i)
And, Mass = Volume×density
M = 4/3πR^3×ρ
or ρ=3M/4πR^3
Using (i),ρ = 3 g/4πRG